The sum of two numbers is 88. Three times the smaller equals twice the larger number,plus 29. Find the numbers.
Answer by Earlsdon(6103)

Let a = the smaller number and b = the larger number, then...
1) a+b = 88 "The sum of two numbers is 88"
2) 3a = 2b+29 "Three times the smaller equals twice the larger plus 29"
Rewrite equation 1) as a = 88-b and substitute into equation 2).
3(88-b) = 2b+29 Simplify and solve for b.
264-3b = 2b+29 Add 3b to both sides.
264 = 5b+29 Subtract 29 from both sides.
235 =5b FInally, divide both sides by 5.
47 = b and a = 88-47 = 41
The two numbers are: 41 and 47